• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 11:20:24 on January22,2004

    Entry number 116420

    This entry is a followup to: 115188

    keyword=Trigger T6 repaired (was broken)

    In the start of the experiment we had intended that trigger T6
    was defined as T3.and.NOT.A1.and.A2
    Mauro discovered that somehow the A2 signal was inefficient.
    Upon access today, Bodo and I found that the timing of the A2
    input was wrong, we fixed it, see figures below.
    The only remaining improvement might be to make the A1 veto
    more conservative by decreasing the gain from 0.5 to 0.3.
    This might get done in the next access, but first let's look
    at the data from this setup.

    Wrong timing of A2 observed today (note, A1 was ok)
    ---------***-------- T3 (22 nsec wide)
    -*******---------- A2 (51 nsec wide, but overlaps only 10% with T3)
    ----************---- A1 (114 nsec wide, 45 nsec earlier than T3)

    Correct timing (had been setup Dec 30, but I suppose the trigger
    database was not saved properly or was corrupted somehow).
    ---------***-------- T3
    -------******-------- A2 (repaired)
    ----*************---- A1 (unchanged)